OpenStudy (anonymous):

CHECK MY ANSWER??? PLEASE Sam stores manure for his farm in a rectangular box. The length of the box is 7 feet, width is 9 inches, and height is 6 inches. Each bag of manure he bought contains 2.625 cubic feet of manure. How many bags of manure will fill the box completely?

OpenStudy (anonymous):

rectangle = length * width * height 7ft = 84in. 84in * 9in = 756in. 756in. * 6in. = 4,536in. 4,536in./2.625 cubic ft. = 1,728in. 1,728in./12in = 144 so he has to buy 144 bags of manure to fill the box. Is this right?

OpenStudy (anonymous):

nvm your answer should be right

OpenStudy (anonymous):

i missed the feet part

OpenStudy (anonymous):

um ok so i am right?

OpenStudy (anonymous):

yeah sorry about the confusion

OpenStudy (anonymous):

Oh ok thanks! (:

OpenStudy (anonymous):

Volume=7×9/12×6/12=21/8=2.625 cubic feet No. of bags=2.625÷2.625=1

OpenStudy (anonymous):

wait what oh goodness im so confused :/

OpenStudy (anonymous):

Volume=7×9/12×6/12=21/8=2.625 cubic feet # of bags=2.625÷2.625.=1....

OpenStudy (anonymous):

yes, i understand that they got that but what i dont understand is why im not right and you guys are. I am not trying to sound rude just confused

OpenStudy (anonymous):

please help? @wolfe8

OpenStudy (anonymous):

2.625 cubic feet=2.625*12*12*12=4536 cubic inch.

OpenStudy (anonymous):

4536/4536=1

OpenStudy (wolfe8):

Sorry I was in class. Do you still need help here?

OpenStudy (anonymous):

it'll be 1 bag of manure you need to convert the length & width in to same unit then width = 0.75ft height = 0.5ft 7*0.5=3sqft is the area volume is 2.625 cubed feet so 1 bag